# On automorphism groups of smooth hypersurfaces

### v1

@@ 4. Fermat-test ratios and proof of most cases
A famous theorem of Jordan states that, for every integer $r\geq 1$, there exists a constant such that any finite subgroup of ${\rm GL}(r,{\mathbb{C}})$ has an abelian normal subgroup of index at most this constant. Consequently, for each integer $r\geq 1$, there are only finitely many finite primitive groups in ${\rm PGL}(r,{\mathbb{C}})$ (up to conjugation), and Collins [Col08a, Theorem A] determined the maximum order ${\rm JC}(r)$ among such groups. Based on canonical bounds and the values of ${\rm JC}(r)$, we introduce the notion of Fermat-test ratio (Definition 4.5) and we show that Fermat-test ratios have many nice properties which will be frequently used in our study. Using Fermat-test ratios, we quickly prove that our main result Theorem 1.1 holds if $n\geq 26$ or $d\geq 18$ (Theorem 4.13).
As a generalization of subdegree sequences of linear groups, the following definition will be used in the definiton of Fermat-test ratios.
@@
Example 4.4.
Let $G$ be as in Example 3.7. Then the subdegree sequence $l(G)$ has the exponential type $2^{1}1^{3}$, and the multiplicity sequence of $l(G)$ is $(1,3)$ which is different from the intrinsic multiplicity sequence $(1,2,1)$ of $G$.
@@
Now we give the definition of Fermat-test ratio which is crucial for our classification of smooth hypersurfaces with large automorphism groups.
@@
By Lemmas 3.6, 3.8 and definition of Fermat-test ratios, we have the following observation.
@@
Let $F=F(x_{1},\ldots,x_{r})$ be a smooth form of degree $d\geq 3$ and $r\geq 1$. Let $\mathcal{H}:=(H_{1},\ldots,H_{s})$ be a primitive-constituent sequence of ${\rm Aut}(F)$ with $r_{1}\geq\cdots\geq r_{m}$, where $r_{i}$ is the degree of $H_{i}$. Then we have
@@
By direct calculation, we find some basic properties of Fermat-test ratios.
@@
The properties of Fermat-test ratios in Lemmas 4.8, 4.9 and 4.11 make computing/bounding values of Fermat-test ratios quite feasible.
@@
    Suppose the subdegrees of $l_{1}$ and $l_{2}$ are disjoint, if $v(l_{2})\geq v(l_{2}^{\prime})$ and $R(l_{2},d)\leq R(l_{2}^{\prime},d)$, then $R(l_{1}+l_{2},d)\leq R(l_{1}+l_{2}^{\prime},d)$; and
@@
The three statements are following from direct calculations. ∎
@@
Let $l$ be a subdegree sequence containing $r_{0}>1$ with multiplicity $k_{0}\geq 1$. Then, for $d\geq 3$, $R(l,d)>R((l+(r_{0}),d))$ holds if one of the following conditions is true:
@@
$\lambda\leq\frac{(r_{0}k_{0})!}{(r_{0}(k_{0}+1))!}\cdot\frac{(k_{0}+1){\rm JC}(r_{0})}{3^{r_{0}-1}}=\frac{(r_{0}k_{0})!}{(r_{0}(k_{0}+1)-1)!}\cdot\frac{{\rm JC}(r_{0})}{r_{0}3^{r_{0}-1}}.$
We notice that the value in the right hand side of the inequality decreases as $k_{0}$ increases. By direct calculation, if $r_{0}=2$, $k_{0}=5$, we have $\lambda\leq\frac{10!\cdot 60}{11!\cdot 2\cdot 3}=\frac{10}{11}<1$. Thus (1) follows. Similarly, one may check if $(r_{0},k_{0})=(3,1)$, $(4,2)$, $(5,1)$, $(6,1)$, $(7,1)$, $(8,1)$, $(9,1)$, $(12,1)$, we have $\lambda<1$.
For $r_{0}=10,11$ or $r_{0}\geq 13$, we have $\lambda<\frac{(r_{0}+1)!}{3^{r_{0}-1}r_{0}^{r_{0}}}<\frac{r_{0}+1}{3^{r_{0}-1}}\cdot\frac{r_{0}^{r_{0}}}{r_{0}^{r_{0}}}<1$. By direct calculation, we find that $R((r_{0}),d)\leq R((r_{0}),3)<1$ for $r_{0}=7$, or $r_{0}\geq 9$. Therefore, $R(r_{0}^{k_{0}},d)<1$ if $r_{0}=7$ or $r_{0}\geq 9$. ∎
@@
$R(3^{1},3)=20/3$, $R(4^{1},3)=40$, $R(4^{2},3)=320/7$, $R(5^{1},3)=8/3$, $R(6^{1},3)=112/3$, $R(8^{1},3)=320/81$.
@@
$l_{1}^{\prime}\text{ is of exponential type}\begin{cases}4^{1}&\quad\text{if }v(l_{1})=4~{}~{}and~{}~{}d=3\\ 2^{a}{~{}~{}with~{}~{}}a=v(l_{1})/2&\quad\text{if }v(l_{1}){~{}~{}is~{}~{}even~{}~{}and~{}~{}}(v(l_{1}),d)\neq(4,3)\\ 2^{a}{~{}~{}with~{}~{}}a=(v(l_{1})-1)/2&\quad\text{if }v(l_{1}){~{}~{}is~{}~{}odd}\\ \end{cases}$
@@
By Lemma 4.9, we have $r_{i}\in\{2,3,4,5,6,8\}$ for all $i$. Then we may assume $l_{1}$ is of exponential type $8^{a_{8}}6^{a_{6}}5^{a_{5}}4^{a_{4}}3^{a_{3}}2^{a_{2}}$ with $a_{i}\geq 0$. Furthermore, as $R(2^{14},3)$, $R(3^{4},3)$, $R(4^{4},3)$, $R(5^{2},3)$, $R(6^{3},3)$, and $R(8^{2},3)$ are all less than 1, applying Lemmas 4.7 and 4.9, we have $a_{8}\leq 1$, $a_{6}\leq 2$, $a_{5}\leq 1$, $a_{4}\leq 3$, $a_{3}\leq 3$, $a_{2}\leq 13$. By calculation, we observe $R(5^{1}3^{1},3)<R(2^{4},3)$, $R(5^{1}3^{3},3)<R(2^{7},3)$, $R(3^{1},3)<R(2^{1},3)$, $R(3^{2},3)<R(2^{3},3)$, $R(3^{3},3)<R(2^{4},3)$, $R(4^{2},3)<R(2^{4},3)$, $R(4^{3},3)<R(2^{6},3)$, $R(5^{1},3)<R(2^{2},3)$, $R(6^{1},3)<R(2^{3},3)$, $R(6^{2},3)<R(2^{6},3)$ and $R(8^{1},3)<R(2^{4},3)$.
If $a_{4}\neq 1$, then by Lemma 4.8 (2)-(3) and replacing subdegrees $3,4,5,6,8$ by $2$ with suitable multiplicities, we have $R(l_{1},d)\leq R(2^{a_{2}^{\prime}},d)$ with $a_{2}^{\prime}=\frac{v(l_{1})}{2}$ (resp. $a_{2}^{\prime}=\frac{v(l_{1})-1}{2}$) when $v(l_{1})$ is even (resp. odd).
If $a_{4}=1$, then similar to the previous case, we have $R(l_{1},d)\leq R(4^{1}2^{b_{2}},d)$ with $b_{2}=\frac{v(l_{1})-4}{2}$ (resp. $b_{2}=\frac{v(l_{1})-5}{2}$) when $v(l_{1})$ is even (resp. odd). Since the ratio $R(4^{1}2^{b_{2}},d)/R(2^{b_{2}+2},d)$ $=\frac{25920}{3600d(b_{2}+1)(b_{2}+2)}$, we find that $R(4^{1}2^{b_{2}},d)\leq R(2^{b_{2}+2},d)$ unless $(b_{2},d)=(0,3)$. Thus, we have $R(l_{1},d)\leq R(2^{b_{2}+2},d)$ unless $(v(l_{1}),d)=(4,3)$.
Moreover, since $R(2^{14},d)<1$ for $d=3$, by Lemmas 4.7 and 4.9, we have $R(l_{1},d)<1$ for all $l_{1}$ satisfying $v(l_{1})\geq 28$. ∎
@@
For later use, here we present (in decreasing order) the integer parts $[R(l_{1}^{\prime},3)]$ of the values $R(l_{1}^{\prime},3)$ for all possible $l_{1}^{\prime}$ with $v(l_{1}^{\prime})\leq 26$ in Lemma 4.11: $[R(2^{5},3)]=105$, $[R(2^{6},3)]=96$, $[R(2^{4},3)]=95$, $[R(2^{7},3)]=74$, $[R(2^{3},3)]=66$, $[R(2^{8},3)]=49$, $R(4^{1},3)=40$, $[R(2^{2},3)]=33$, $[R(2^{9},3)]=29$, $[R(2^{10},3)]=15$, $R(2^{1},3)=10$, $[R(2^{11},3)]=7$, $[R(2^{12},3)]=3$, $[R(2^{13},3)]=1$.
@@
Let $F=F(x_{1},\dots,x_{n+2})$ be a smooth form of degree $d$, where $n\geq 1$, $d\geq 3$. Suppose $X_{F}$ is not isomorphic to the Fermat hypersurface in ${\mathbb{P}}^{n+1}$ of degree $d$. If $n\geq 26$ or $d\geq 18$, then $|{\rm Aut}(X_{F})|<d^{n+1}\cdot(n+2)!$.
@@
Suppose the subdegree sequence $l:=l({\rm Aut}(F))$ of ${\rm Aut}(F)$ is of exponential type $r_{1}^{k_{1}}\cdots r_{m}^{k_{m}}$ with $m\geq 1$ and $k_{i}>0$ for all $i$. By Theorem 3.9, we may assume $r_{1}>1$. Then by Lemma 4.6, it suffices to show $R(l,d)<1$. Let $\mathcal{A}=\{i|\;r_{i}>1,R(r_{i}^{k_{i}},d)\geq 1,1\leq i\leq m\}.$ We define $l_{1}$ to be the sum of all the subdegree sequences of exponential types $r_{j}^{k_{j}}$ ($j\in\mathcal{A}$). Let $l_{2}:=l-l_{1}$.
Case $n\geq 26$. If $l=l_{1}$ or $l=l_{2}$, then by Lemmas 4.11 and 4.8, $R(l,d)<1$. Then we may assume that $l\neq l_{1},l_{2}$ and $v(l_{1})<28$. Since $n+2=v(l_{1}+l_{2})\geq 28$, by Lemmas 4.8, 4.11 and Remark 4.12, we have
@@
Case $d\geq 18$. By Lemma 4.7, it suffices to show $R(l,18)<1$. For positive integers $r_{0}>1$ and $k_{0}>0$, by Lemma 4.9 and direct calculation, $R(r_{0}^{k_{0}},18)\geq 1$ if and only if $(r_{0},k_{0})=(2,1)$. Thus, by Lemma 4.11, either $l_{1}=l_{1}^{\prime}$ of exponential type $2^{1}$ or $l=l_{2}$, which implies $R(l,18)<1$ by Lemma 4.8 again. This completes the proof of the theorem. ∎
By similar arguments in the proof of Theorem 4.13, we obtain additional properties of Fermat-test ratios, which will be used in the subsequent sections.
@@
    If $l$ has at least two distinct subdegrees, then $R(l,d)<60$.

### v2

@@ 4. Fermat-test ratios and proof of most cases
A famous theorem of Jordan [Jor78] states that, for every integer $r\geq 1$, there exists a constant such that any finite subgroup of ${\rm GL}(r,{\mathbb{C}})$ has an abelian normal subgroup of index at most this constant. Consequently, for each integer $r\geq 1$, there are only finitely many finite primitive groups in ${\rm PGL}(r,{\mathbb{C}})$ (up to conjugation), and Collins [Col08a, Theorem A] determined the maximum order ${\rm JC}(r)$ among such groups. Note that ${\rm JC}(r)={\rm J}(r)$ for $r\geq 71$, but ${\rm JC}(r)$ and ${\rm J}(r)$ are often different for $r<71$ (see [Col07, Theorems A, B and D]). In §4.1, based on canonical bounds and the values of ${\rm JC}(r)$, we introduce the notion of Fermat-test ratio (Definition 4.5) and we show that Fermat-test ratios have many nice properties which will be frequently used in our study. In §4.2, using Fermat-test ratios, we quickly prove that our main result Theorem 1.1 holds if $n\geq 26$ or $d\geq 18$ (Theorem 4.14).
4.1. Fermat-test ratios
As a generalization of subdegree sequences of linear groups, the following concepts will be used in the definition of Fermat-test ratios.
@@ 4.1. Fermat-test ratios
In practice, we often use $\mathcal{H}$ to denote a primitive-constituent sequence.
Example 4.2.
Let $\widetilde{G}$ be as in Example 3.8. Then the subdegree sequence $l(\widetilde{G})=(2,1,1,1)$ has the exponential type $2^{1}1^{3}$, and the multiplicity sequence of $l(\widetilde{G})$ is $(1,3)$ which is different from the intrinsic multiplicity sequence $(1,2,1)$ of $\widetilde{G}$. The total degree $v(l(\widetilde{G}))$ is 5.
@@
Now we give the definition of Fermat-test ratios which is crucial for our classification of smooth hypersurfaces with large automorphism groups.
@@
The following simple fact is useful for our study.
@@
By Lemmas 3.7, 4.6 and definition of Fermat-test ratios, we have the following observation.
Lemma 4.7.
Let $F=F(x_{1},\ldots,x_{r})$ be a smooth form of degree $d\geq 3$ and $r\geq 1$. Let $\mathcal{H}:=(H_{1},\ldots,H_{s})$ be a primitive-constituent sequence of ${\rm Aut}(F)$ with $r_{1}\geq\cdots\geq r_{s}$, where $r_{i}$ is the degree of $H_{i}$. Then we have
@@
By Definition 4.5, we find some basic properties of Fermat-test ratios.
@@
The properties of Fermat-test ratios in Lemmas 4.9, 4.10 and 4.12 make computing/bounding values of Fermat-test ratios quite feasible.
@@
    Suppose the subdegrees of $l_{1}$ and $l_{2}$ are disjoint. Let $l_{2}^{\prime}$ be a subdegree sequence. If $v(l_{2})\geq v(l_{2}^{\prime})$ and $R(l_{2},d)\leq R(l_{2}^{\prime},d)$, then $R(l_{1}+l_{2},d)\leq R(l_{1}+l_{2}^{\prime},d)$; and
@@
We only give the proof of case (1) and the other cases are similar. Suppose the exponential type of $l:=l_{1}+l_{2}$ is $r_{1}^{k_{1}}r_{2}^{k_{2}}\cdots r_{m}^{k_{m}}$. Then we may write $l_{1}=r_{1}^{k^{\prime}_{1}}r_{2}^{k^{\prime}_{2}}\cdots r_{m}^{k^{\prime}_{m}}$ and $l_{2}=r_{1}^{k^{\prime\prime}_{1}}r_{2}^{k^{\prime\prime}_{2}}\cdots r_{m}^{k^{\prime\prime}_{m}},$ where $k^{\prime}_{i}\geq 0$, $k^{\prime\prime}_{i}\geq 0$ and $k_{i}=k^{\prime}_{i}+k^{\prime\prime}_{i}>0$. From Lemma 4.6, we conclude that $k_{i}!\geq k^{\prime}_{i}!\cdot k^{\prime\prime}_{i}!$ and the equality holds if and only if $k^{\prime}_{i}\cdot k^{\prime\prime}_{i}=0$. By Definition 4.5, we have
$R(l,d)=R(l_{1},d)R(l_{2},d)\cdot\frac{v(l_{1})!\cdot v(l_{2})!}{v(l_{1}+l_{2})!}\cdot\prod_{i=1}^{m}\frac{k_{i}!}{k_{i}^{\prime}!\cdot k_{i}^{\prime\prime}!}\geq R(l_{1},d)R(l_{2},d)\cdot\frac{1}{\tbinom{v({l_{1}+l_{2}})}{v({l_{1}})}},$
and the equality holds if and only if $k^{\prime}_{i}\cdot k^{\prime\prime}_{i}=0$ for all $i$ (i.e., the subdegrees of $l_{1}$ and $l_{2}$ are disjoint). ∎
@@
Let $l$ be a subdegree sequence containing $r_{0}>1$ with multiplicity $k_{0}\geq 1$. Then, for $d\geq 3$, $R(l,d)>R(l+(r_{0}),d)$ holds if one of the following conditions is true:
@@
$\lambda\leq\frac{(r_{0}k_{0})!}{(r_{0}(k_{0}+1))!}\cdot\frac{(k_{0}+1){\rm JC}(r_{0})}{3^{r_{0}-1}}=\frac{(r_{0}k_{0})!}{(r_{0}(k_{0}+1)-1)!}\cdot\frac{{\rm JC}(r_{0})}{r_{0}3^{r_{0}-1}}.$ (4.2)
We notice that the value in the right hand side of the inequality decreases as $k_{0}$ increases. By (4.2), if $r_{0}=2$, $k_{0}=5$, we have $\lambda\leq\frac{10!\cdot 60}{11!\cdot 2\cdot 3}=\frac{10}{11}<1$. Thus (1) follows. Similarly, one may check if $(r_{0},k_{0})=(3,1)$, $(4,2)$, $(5,1)$, $(6,1)$, $(7,1)$, $(8,1)$, $(9,1)$, $(12,1)$, we have $\lambda<1$.
For $r_{0}=10,11$ or $r_{0}\geq 13$, we have $\lambda<\frac{(r_{0}+1)!}{3^{r_{0}-1}r_{0}^{r_{0}}}<\frac{r_{0}+1}{3^{r_{0}-1}}\cdot\frac{r_{0}^{r_{0}}}{r_{0}^{r_{0}}}<1$. By Lemma 4.8 and Definition 4.5, we find that $R((r_{0}),d)\leq R((r_{0}),3)<1$ for $r_{0}=7$, or $r_{0}\geq 9$. Therefore, $R(r_{0}^{k_{0}},d)<1$ if $r_{0}=7$ or $r_{0}\geq 9$. ∎
@@
For the proof of Lemma 4.12, we record the values of some Fermat-test ratios. $R(2^{13},3)=16000000000/12649365729$, $R(3^{1},3)=20/3$, $R(3^{3},3)=200/189$, $R(4^{1},3)=40$, $R(4^{2},3)=320/7$, $R(4^{3},3)=2560/231$, $R(5^{1},3)=8/3$, $R(6^{1},3)=112/3$, $R(6^{2},3)=896/297$, $R(8^{1},3)=320/81$.
@@
$l_{1}^{\prime}\text{ is of exponential type}\begin{cases}4^{1}&\quad\text{if }v(l_{1})=4~{}~{}and~{}~{}d=3\\ 2^{a}{~{}~{}with~{}~{}}a=v(l_{1})/2&\quad\text{if }v(l_{1}){~{}~{}is~{}~{}even~{}~{}and~{}~{}}(v(l_{1}),d)\neq(4,3)\\ 2^{a}{~{}~{}with~{}~{}}a=(v(l_{1})-1)/2&\quad\text{if }v(l_{1}){~{}~{}is~{}~{}odd.}\\ \end{cases}$
@@
By Lemma 4.10, we have $r_{i}\in\{2,3,4,5,6,8\}$ for all $i$. Then we may assume $l_{1}$ is of exponential type $8^{a_{8}}6^{a_{6}}5^{a_{5}}4^{a_{4}}3^{a_{3}}2^{a_{2}}$ with $a_{i}\geq 0$. Furthermore, as $R(2^{14},3)$, $R(3^{4},3)$, $R(4^{4},3)$, $R(5^{2},3)$, $R(6^{3},3)$, and $R(8^{2},3)$ are all less than 1, applying Lemmas 4.8 and 4.10, we have $a_{8}\leq 1$, $a_{6}\leq 2$, $a_{5}\leq 1$, $a_{4}\leq 3$, $a_{3}\leq 3$, $a_{2}\leq 13$. By calculations, we observe $R(5^{1}3^{1},3)<R(2^{4},3)$, $R(5^{1}3^{3},3)<R(2^{7},3)$, $R(3^{1},3)<R(2^{1},3)$, $R(3^{2},3)<R(2^{3},3)$, $R(3^{3},3)<R(2^{4},3)$, $R(4^{2},3)<R(2^{4},3)$, $R(4^{3},3)<R(2^{6},3)$, $R(5^{1},3)<R(2^{2},3)$, $R(6^{1},3)<R(2^{3},3)$, $R(6^{2},3)<R(2^{6},3)$ and $R(8^{1},3)<R(2^{4},3)$.
[…]
Moreover, since $R(2^{14},d)<1$ for $d=3$, by Lemmas 4.8 and 4.10, we have $R(l_{1},d)<1$ for all $l_{1}$ satisfying $v(l_{1})\geq 28$. ∎
@@
For later use, here we present (in decreasing order) the values $R(l_{1}^{\prime},3)$ for all possible $l_{1}^{\prime}$ with $v(l_{1}^{\prime})\leq 26$ in Lemma 4.12: $R(2^{5},3)=20000/189$, $R(2^{6},3)=200000/2079$, $R(2^{4},3)=2000/21$, $R(2^{7},3)=2000000/27027$, $R(2^{3},3)=200/3$, $R(2^{8},3)=4000000/81081$, $R(4^{1},3)=40$, $R(2^{2},3)=100/3$, $R(2^{9},3)=40000000/1378377$, $R(2^{10},3)=400000000/26189163$, $R(2^{1},3)=10$, $R(2^{11},3)=4000000000/549972423$, $R(2^{12},3)=40000000000/12649365729$, $R(2^{13},3)=16000000000/12649365729$.
4.2. Proof of most cases
@@
Let $F=F(x_{1},\dots,x_{n+2})$ be a smooth form of degree $d$, where $n\geq 1$, $d\geq 3$. Suppose that $X_{F}$ is not isomorphic to the Fermat hypersurface in ${\mathbb{P}}^{n+1}$ of degree $d$. If $n\geq 26$ or $d\geq 18$, then $|{\rm Aut}(X_{F})|<d^{n+1}\cdot(n+2)!$.
@@
Suppose the subdegree sequence $l:=l({\rm Aut}(F))$ of ${\rm Aut}(F)$ is of exponential type $r_{1}^{k_{1}}\cdots r_{m}^{k_{m}}$ with $m\geq 1$ and $k_{i}>0$ for all $i$. By Theorem 3.9, we may assume $r_{1}>1$. Then by Lemma 4.7, it suffices to show $R(l,d)<1$. Let $\mathcal{A}:=\{i|\;r_{i}>1,R(r_{i}^{k_{i}},d)\geq 1,1\leq i\leq m\}.$ We define $l_{1}$ to be the sum of all the subdegree sequences of exponential types $r_{j}^{k_{j}}$ ($j\in\mathcal{A}$). Let $l_{2}:=l-l_{1}$.
Case $n\geq 26$. If $l=l_{1}$ or $l=l_{2}$, then by Lemmas 4.12 and 4.9 (1), $R(l,d)<1$. Then we may assume that $l\neq l_{1},l_{2}$ and $v(l_{1})<28$. Since $n+2=v(l_{1}+l_{2})\geq 28$, by Lemmas 4.9 (1), 4.12 and Remark 4.13, we have
@@
Case $d\geq 18$. By Lemma 4.8, it suffices to show $R(l,18)<1$. For positive integers $r_{0}>1$ and $k_{0}>0$, from Lemma 4.10, we conclude that $R(r_{0}^{k_{0}},18)\geq 1$ if and only if $(r_{0},k_{0})=(2,1)$. Thus, by Lemma 4.12, either $l_{1}=l_{1}^{\prime}$ of exponential type $2^{1}$ or $l=l_{2}$, which implies $R(l,18)<1$ by Lemma 4.9 (1) again (in fact, if $l=l^{\prime}=2^{1}$, then $R(l,18)=\frac{R(2^{1},18)\cdot R(l_{2},18)}{\tbinom{n+2}{2}}\leq\frac{R(2^{1},18)}{\tbinom{n+2}{2}}<1$; if $l=l_{2}$, then $R(l,18)\leq\prod_{1\leq i\leq m}R(r_{i}^{k_{i}},18)<1$ since $R(r_{i}^{k_{i}},18)<1$ for all $r_{i}>1$). This completes the proof of the theorem. ∎
By similar arguments in the proof of Theorem 4.14 and running some local computations, we obtain additional properties of Fermat-test ratios, which will be used in the subsequent sections.
@@
    If $l$ has at least two distinct subdegrees, then $R(l,d)<60$;
@@
To get the bounds in (1), (2), (3), it suffices to consider the case $d=3$ by Lemma 4.8. Note that 106 is a uniform bound for Fermat-test ratios.